A rectangular patio is 7 ft longer than it is wide. Determine the dimensions of the patio if it measures 13 ft diagonally.
Answer by checkley75(3666)

x^2+(x+7)^2=13^2
x^2+x^2+14x+49=169
2x^2+14x+49-169=0
2x^2+14x-120=0
2(x^2+7x-60)=0
2(x-5)(x+12)=0
x-5=0
x=5 ans. for the length.
x+7=12 ans. for the width.
Proof:
5^2+12^2=13^2
25+144=169
169=169